OpenStudy (roku):

another mole question show the working for me please i am not sure how to do this problem the eqn. for photosynthesis.. if 100g of water react with excess oxygen, how many liters of oxygen will be produced? assume stp

OpenStudy (roku):

i got 124.5 dm^3

OpenStudy (anonymous):

Bio isn't exactly my best subject :/ can you show me the working for photosynthesis? I'll be able to help you then.

OpenStudy (roku):

6CO2carbon dioxide + 6H2O water +light-> C6H12O6 glucose+ 6(O2) oxygen

OpenStudy (roku):

the mole ratio would be re written as 1:1 right?

OpenStudy (anonymous):

18 * 6 gm of h20 reacr to form 32 * 6 gm o2 then....100gm of water react to form hw many gm of o2?

OpenStudy (anonymous):

6CO2+ 6H2O -> C6H12O6 + 6O2 Mol H2O = 100/18 = 5.56 6 Mol H2O -> 6 Mol O2 5.56 mol H2O -> 1* 5.56 mol O2 = 5.56 mol O2 5.56*22.4= 124.544 dm^3

OpenStudy (anonymous):

@Roku : yeah, the mol ratio would be written as that when you compare (:

OpenStudy (anonymous):

32 * 6 / 18 * 6 * 100 = 177.78 gm.

OpenStudy (anonymous):

Then you just convert it into litre! I forgot about that :/

OpenStudy (anonymous):

22.4 L = 1 mole Moles = 177.78 /32 = 5.5 So........ 22.4 * 5.55 = 124.544 L

OpenStudy (roku):

thanks guys

OpenStudy (anonymous):

Welcome....)

OpenStudy (roku):

yahoo i dont really understand your way completely

OpenStudy (roku):

so it would be 124544cm^3 ??

OpenStudy (anonymous):

^ yeah.